Electronic Journal: Southwest Journal of Pure and Applied Mathematics Internet: http://rattler.cameron.edu/swjpam.html ISSN 1083-0464 Vol. No. 01, 1995, pp. 37-45. Submitted: June 8, 1995. Published: October 26, 1995.

A PROBABILISTIC APPROACH TO THE ESTIMATION OF THE DISTRIBUTION

OF THE TARGET LOCATION POINT

M.A. Tabatabai

Abstract. A probabilistic approach is used to estimate the probability distribution of the target location point. The results can be used in missile-testing and target acquisition fire support model simulation. The target may be a mobile multiple launcher rocket, an antiaircraft gun or a ship in mid-ocean. In each case, there is a time lag between the detection of the target and the arrival of the projectile.

(1991) A.M.S. (MOS) Subject Classification Codes. 60, 62.

Key Words and Phrases. Probability density function, Gamma function.

Department of Mathematical Sciences, Cameron University, Lawton, OK 73505 E-mail Address: mohammad@tiger.cameron.edu \begin{matrix}c \\ ◯ \end{matrix}1995 Cameron University

Typeset by AMS-TEX